OpenStudy (anonymous):

Biologists have observed that the chirping rate of crickets of a certain species is related to temperature, and the relationship appears to be very nearly linear. A cricket produces 120 chirps per minute at 70°F and 168 chirps per minute at 80°F. (a) Find the linear equation that relates the temperature t and the number of chirps per minute n. (b) If the crickets are chirping at 138 chirps per minute, estimate the temperature. (Round your answer to the nearest degree.)

zepdrix (zepdrix):

Hmm based on the way they worded part b, it sounds like they want you to write this equation as a function of Temperature --> n(t) So we need to find a linear equation to model this. It will be of the form n(t)=mt+b to find m: \[m=\frac{ y_2 - y_1 }{ x_2 - x_1 }\]\[m=\frac{ n_2 - n_1 }{ t_2 - t_1 }\] Let's get our data into ordered pairs and find the slope.

zepdrix (zepdrix):

mm i did that backwards, i meant a function of chirps lol... crap :D

OpenStudy (anonymous):

10/48= 5/24

zepdrix (zepdrix):

it'll stil work out fine though :)

zepdrix (zepdrix):

did you put chirps or temperature on top? :o

zepdrix (zepdrix):

you put temperature, thats fine, lets just rewrite our formula in terms of chirps then :) \[t(n)=mn+b\]

zepdrix (zepdrix):

so we found our m, the slope of our function. m=5/24 \[t(n)=\frac{ 5 }{ 24 }n+b\] plug in one of your coordinate pairs to find a b value :D

OpenStudy (anonymous):

168=5/24n+80

zepdrix (zepdrix):

woops, you plugged 80 in for b. our ordered pair is of the form (n,t)

OpenStudy (anonymous):

im confused please show me

zepdrix (zepdrix):

t(n)=(5/24)n+b plugging in (n,t(n)) which is (168, 80) -> 80=(5/24)(168)+b

OpenStudy (anonymous):

ok?

zepdrix (zepdrix):

solve for b :)

OpenStudy (anonymous):

b=45

zepdrix (zepdrix):

yay good! :) ok we have an equation for the line \[t(n)=(5/24)n+45\] now lets use that equation to find the temperature when the number of chirps per minute is 138. \[t(138)=(5/24)(138)+45\]

OpenStudy (anonymous):

do i have to put t(n) or can i just put n

zepdrix (zepdrix):

? :o

zepdrix (zepdrix):

t(n) is our function output, its similar to f(x), it's our y variable. It's not the same as n, if that's what you were asking <:o

OpenStudy (anonymous):

what about just putting t instead of t(n) im trying to enter it on an online hw thing

zepdrix (zepdrix):

oh i see, yah try t=(5/24)n+45

OpenStudy (anonymous):

Thanks :)

zepdrix (zepdrix):

yay gj, were you able to get an answer for b also?

OpenStudy (anonymous):

yes :)